### v1

# Long Context Scaling: Divide and Conquer via Multi-Agent Question-driven Collaboration

Abstract
Processing long contexts has become a critical capability for modern large language models (LLMs). Existing works leverage agent-based divide-and-conquer methods for processing long contexts. But these methods face crucial limitations, including prohibitive accumulated latency and amplified information loss from excessive agent invocations, and the disruption of inherent textual dependencies by immoderate partitioning. In this paper, we propose a novel multi-agent framework XpandA (Expand-Agent) coupled with question-driven workflow and dynamic partitioning for robust long-context processing. XpandA overcomes these limitations through: 1) dynamic partitioning of long texts, which adaptively modulates the filling rate of context windows for input sequences of vastly varying lengths; 2) question-guided protocol to update flat information ensembles within centralized shared memory, constructing consistent inter-agent knowledge across partitions; and 3) selectively replaying specific partitions based on the state-tracking of question-information couples to promote the resolution of inverted-order structures across partitions (e.g., flashbacks). We perform a comprehensive evaluation of XpandA on multiple long-context benchmarks with length varying from 1k to 1M, demonstrating XpandA’s feasibility for processing ultra-long sequences and its significant effectiveness in enhancing the long-context capabilities of various LLMs by achieving 20% improvements and 1.5x inference speedup over baselines of full-context, RAG and previous agent-based methods.
1 Introduction
Like human cognition—which depends on long-term memory, accumulated knowledge, and life experience—Artificial General Intelligence (AGI) development requires the ability to associate distant information and retain context over extended content [38; 58]. Long-context tasks, such as question answering [76; 92; 68], summarization [105] and coding [7], consistently present significant challenges due to issues like hallucination and overlooking critical information [50].
Recent efforts to tackle these challenges have focused on three key directions: training data strategy, architecture and workflow design. The training data strategy conducts long-context data filtering [86; 71], mixture [23; 33], and synthesis [24; 27] on pre-training and post-training phase. Architecture methods employ position embeddings [73; 64; 25; 74] and attention mechanisms such as transformer-based [5; 88; 108; 95; 22], linear-complexity [65] and hybrid architectures [59; 47; 53]. Despite these efforts, the lack of extensibility in context windows leads to input truncation when processing longer sequences, while training long-context language models entails prohibitively high computational costs[5]. Workflow design, including external memory modules [80; 56], retrieval-augmented generation (RAG) [8; 34], and agent-based methods (LongAgent [101], Chain of Agents [100]), etc., can mitigate the challenges mentioned above but introduce issues such as information loss for inappropriate extraction [100], Inter-Agent Misalignment and Specification Violation inherent in the multi-agent system [9]. Furthermore, the most concerning aspect of agent-based methods is the frequent model invocation. This not only potentially amplifies intermediate errors but also leads to an accumulation of model call latency.
Motivated by the aforementioned challenges, we propose a multi-agent framework XpandA (Expand-Agent) that incorporates fine-grained question-driven communication protocols and, for the first time, extends the input sequence length of agent-based methods beyond 128k tokens. By introducing a Saturating Function within its partitioning strategy, XpandA can practically process input sequences with vastly varying lengths and achieve a trade-off between unit inference time and accumulated model call latency. Furthermore, XpandA effectively guides the model’s attention transfer across partitions via its question-driven communication protocol, thereby mitigating information dependency fragmentation caused by partitioning. Inspired by Activity On Vertex (AOV) networks, we implement an iterative selective replay mechanism based on collected information-question couples. This approach effectively resolves potential topological structures of information across different text partitions (such as narratives with flashbacks or interpolations that cannot be resolved sequentially).
We conduct comprehensive experiments on long-context benchmarks with diverse tasks of vastly varying lengths (from 1k to 1M). Utilizing open-source LLMs, we compare XpandA with strong baselines of full-context, Retrieval-Augmented Generation (RAG), and prior Multi-agent methods, demonstrating its significant improvement over all baselines by up to 20% and 1.5x inference speedup. In summary, our contributions to the research community of long-context processing include:
  1. 1.
    Proposing a multi-agent framework with fine-grained question-driven communication protocols that can practically expand the LLM’s context window beyond 128k without training.
  2. 2.
    Developing a dynamic partitioning strategy that achieves asymptotically optimal inference speedup while enhancing inference performance by trade-off between unit inference time and accumulated model call latency with input sequences of vastly varying lengths (from 1k to 1M).
  3. 3.
    Establishing a paradigm for multi-agent system in long-context tasks domain to mitigate inter-agent misalignment and specification violations using principled prompt and structured workflow.
Figure 1: An Overview of XpandA. A plug-and-play, robust, interpretable, and efficient multi-agent system framework that introduces a dynamic chunking strategy (a) and selective replay (b) in its question-driven workflow (c). XpandA accepts long context input and splits the input into chunks with overlap, and each chunk is fed to corresponding Explorer Agents $\{E[i,1]\ |\ i\in{1,2,3}\}$ sequentially. Then Decider D1 decides the action of the next step is to Replay or Conclude. If Replay is proposed, $E[i,2]$ will be called in a reverse direction with a start point according to the question state tracker.
2 Related Work
2.1 Long Context Language Modeling
Long context processing in LLMs has recently become a central focus [49], driven by tasks such as long-document QA [35; 15; 76; 55; 21], summarization [29; 105], retrieval-augmented generation (RAG) [39], in-context learning [46; 90], and large-codebase understanding [32; 7]. These require models to locate, aggregate, and reason over dispersed information, while overcoming "lost in the middle"[50] issues. Existing approaches fall into three main categories: Training data strategies, architectural and workflow design [48]. Data strategies enhance a model’s capacity to process lengthy texts through the filter [86; 71], mixture [23; 33], and synthesis [24; 27] of long-context data in both pre-training and post-training phases. Architecture design aims to develop model structures capable of effectively handling extended texts by position embeddings [73; 64; 25; 74] and attention mechanisms such as transformer-based [5; 88; 108; 95; 22], linear-complexity [65] and hybrid architectures [59; 47; 53]. Workflow-based methods minimize model changes, using techniques like external memory [80; 56], RAG [8; 34], or agent-based reasoning [11; 45; 100], to further improve the long context capability.
2.2 LLM Reasoning with Search and Planning
Human problem-solving relies on decomposing complex tasks into manageable steps [19], a paradigm successfully adapted to LLMs. Early work by [85] and [36] introduced Chain-of-Thought (CoT) prompting, using intermediate steps to guide reasoning, while later studies like [20] and [106] proposed iterative decomposition into subtasks. To enhance reasoning structures, [107] and [3] developed frameworks for self-discovering and dynamically adapting reasoning modules. Further, methods like [17] and [99] address error correction in intermediate steps, particularly for smaller models. Beyond linear reasoning, exploration-based approaches leverage search over multiple paths, inspired by human problem-solving [72]. While early methods like self-consistency [81] generated limited diversity, advanced frameworks such as Tree-of-Thoughts [93] and Graph-of-Thoughts [6] enable finer-grained branching. Aggregation strategies further improve robustness, ranging from ensemble-based voting [77; 44] to verifier-guided selection [82].
2.3 Multi-Agent LLMs
Research efforts in Multi-Agent Systems have encompassed diverse aspects of collaboration, covering mechanisms like cooperation [14; 70], competition [12; 102], and coopetition [1; 16], as well as strategies such as rule-based [97; 109], role-based [14; 28], and model-based approaches [42; 89]. Furthermore, other key considerations include communication structures (centralized [31; 61], decentralized [14; 94], and hierarchical [41; 52]), which offer scenario-dependent advantages and disadvantages, and coordination architectures (static or dynamic [13; 30; 83]), providing trade-offs between stability and flexibility. MASs demonstrate broad potential for applications across numerous domains, including 5G/6G networks [84; 104], Natural Language Generation (NLG) and Question Answering (QA) [31; 75; 87] and socio-cultural simulations [40; 60], among others. Nevertheless, the field still confronts numerous challenges [10]. These include achieving unified governance, understanding the inherent limitations of individual agents [97; 69], scalability, efficient resource management [79], developing robust evaluation benchmarks [51; 63] and addressing ethical risks and safety concerns [2; 18; 57].
Input: Context $C$, Query $Q$.
Parameters: LLM$(\theta)$, instruct prompts $(I_{E},I_{D})$, chunk size $K$, chunk overlap $P$, max replay times MRT
Output: Answer $A$.
Split $C$ into $x$ chunks $\{c_{1},c_{2},\cdots,c_{x}\}$;
Create empty dictionary $T$, $P$;
Initialize $o\leftarrow 1$, $rev\leftarrow 1$;
[…]
       for $i$ in $o,o+rev,\cdots$ do
             $T_{i},P_{i}\leftarrow\text{Explorer}(I_{E},Q,T,P,c_{i})$;
            $T\leftarrow\text{Merge}(T,T_{i})$;
             $P\leftarrow\text{Merge}(P,P_{i})$;
            for every $k\in P_{i}.\text{keys}$ and $k\in T.\text{keys}$ do
                   $\text{Delete}(k,T[k])$ from $T$;
             end for
       end for
      $\text{Action},\hat{A}\leftarrow\text{Decider}(I_{D},Q,P)$;
       if Replay in Action then
             $o\leftarrow\begin{cases}\max(\min(T.values-1,1)&\text{if }rev>0\\ \min(\max(T.values+1,l)&\text{otherwise}\end{cases}$
            $rev\leftarrow-rev$;
[…]
       end if
      else Break;
end while
$A\leftarrow\text{Parse}(\hat{A})$;
return $A$;
Algorithm 1 XpandA Workflow
3 Method
Figure 1 illustrates our XpandA framework, which operates in three stages. In Stage 1, Dynamic partition strategy is employed to decompose the input sequence of vastly varying length into chunks of appropriate length. In Stage 2, text chunks are assigned one-to-one to Explorers. Explorers then process the chunks sequentially by: decomposing problems, searching information and then generating new problems. Problems and relevant information are added to shared information memory. In Stage 3, Decider evaluates information completeness: if sufficient, it generates the final answer; otherwise, it proposes a selective replay trajectory of potentially informative chunks.
3.1 Stage 1: Dynamic Partition of Long Input Sequence
The Dynamic Partition method uses a saturation function: for shorter inputs, it increases unit chunk length; for longer inputs, the length saturates and chunk count grows. Let $w$ be the input token length, $n$ the target chunk count, $[L,K]$ the range of overlap, and $M$ the maximum chunk size. The partitioning follows:
Chunk $\text{Chunk size}=\begin{cases}\left\lceil\frac{w}{n}\right\rceil&\text{if }\frac{w}{n}\leq M\\ M&\text{if }\frac{w}{n}>M\end{cases}$ (1)
[…]
Number of $\text{Number of chunks}=\left\lceil\frac{w}{M-\delta}\right\rceil$ (2)
where $\delta$ (with increase rate $\alpha$, upper limit $K$ and lower limit $L$)  [^1] The chunking strategy in Xpanda sets $n=3,L=10,K=2000,\alpha=0.1,M=102400$. equals to:
$\delta=max(L,min(\alpha*w,K))$ (3)
3.2 Stage 2: Question-guided Exploration
In Stage 1: XpandA operates on a sequence of $l$ Explorers. Each Explorer accepts the concatenation of a document chunk $c_{i}$, user query $q$, instruction for a specific reasoning task $I_{D}$ and gathered information, denoted as "(question,[answer]) pairs" $P$. Explorer is instructed to output enlarged $P_{i}$ and generate newly-generated problems included in "Unsolved Problem Tracer" $T_{i}$ expressed as:
$T_{i},P_{i}\leftarrow\text{Explorer}(I_{E},Q,T,P,c_{i})$ (4)
$T$,$P$ will be updated with $T_{i}$,$P_{i}$ when the processed chunk is switched.
$T\leftarrow Merge(T,T_{i})$ (5)
$P\leftarrow Merge(P,P_{i})$
In the workflow, intricate problems are broken down into sub-problems $q_{x}^{y}$ (the $y^{th}$ problem from $x^{th}$ chunk), which are supplemented by new information, and then new problems are proposed based on the newly-acquired information $a_{n}^{m}$ (the $m^{th}$ answer for the $n^{th}$ problem). The process can be specified by:
$q_{i}^{n}\leftarrow Expand(P_{i})$ (6)
$a_{n}^{j}\leftarrow Answer(q_{x}^{n}),\forall x<=i$
$Expand$ function has two types of behaviors: 1) Breaking down the generalized problem into sub-problems, and 2) expanding new problem entities based on a problem and the corresponding information. While the loop iterates, the problem-solving focus of large language model is gradually expanded across the chunks.
3.3 Stage 3: Iterative Selective Replay
Chunking inherently suffers from information loss: interrelated information across different chunks may be lost if not added to public memory, meaning the model cannot spontaneously notice dangling nodes in broken reasoning chains beyond its context window. For instance, if reasoning about information in chunk a depends on reasoning about information in chunk b (where $b<a$), the model cannot initiate reasoning for chunk b. To address this, we propose a heuristic replay strategy. The Unsolved Problem Tracer $T$ established in the previous stage is designed precisely for this strategy. Upon completion of each unidirectional reasoning phase, the Decider is activated to assess whether the currently gathered information suffices for reasoning and determines whether to initiate replay.
$Action,\hat{A}\leftarrow\text{Decider}(I_{D},Q,P)$ (7)
The replay performs backward search from the last unsolved chunk, applying two heuristics: 1) the target chunk cannot answer all pending questions, and 2) all relevant information precedes it:
$o\leftarrow\begin{cases}\max(\min(T.values-1,1)&\text{if }rev>0\\ \min(\max(T.values+1,l)&\text{otherwise}\end{cases}$ (8)
This dependency forms an AOV network where chunks are vertices and dependencies are edges. When MRT (max replay times) $>x-1$, XpandA guarantees to complete dependency resolution A if LLM can finish information extraction in a chunk unit.
4 Experiment
4.1 Experiment Setup
[…]
  • Question Answering. HotpotQA [92] is a Wikipedia-based multihop QA dataset. It requires reasoning across multiple passages to find the answer. NarrativeQA [68] is a dataset focusing on answering questions based on stories or scripts, including understanding of important elements such as characters, plots, themes, etc. loogle-MR-mixup and loogle-CR-mixup originate from LooGLE [43] Long-dependency QA task, specifically the Multiple information Retrieval and Comprehension and Reasoning subtasks and are mixed up with mix up distracting documents and supporting documents of stepwise growing length.
  • Retrieval. OpenAI-MRCR is a long context dataset for benchmarking an LLM’s ability to distinguish between multiple needles hidden in context. The evaluation consists of retrieval of the response corresponding to a specific instance from multi-turn synthetic dialogue.
[…]
  • Misc. We pick GraphWalks. Graphwalks is designed to require reasoning across multiple positions in the context and cannot be solved sequentially. In Graphwalks, the model is given a graph represented by its edge list and asked to perform an operation to search the parent or the breadth-first search (BFS) results from certain nodes.
LLM Backbones. The backbone models we deploy through include the well-established and popular open-source model Llama3.1-8B-Instuct [26], Qwen2.5-7B-Instruct [66] as well as the latest long-context language model Qwen2.5-7B-Instruct-1M [91] model. We choose efficient engine vllm [37] and 4*A100_80G GPU as infrastructures for the experiments to handle immense GPU memory required for ultra-long input sequences up to 1M tokens.
Baselines. Full-context directly places the query and the long text into the context window where truncation may occur. RAG is implemented through RAG-Fusion [67] by utilizing reciprocal rank fusion of relevance and word frequency ranking to reach strong and robust retrieval performance. We introduce previous Agent-based works CoA [100] and LongAgent [101] to enhance the comprehensiveness of our baseline to highlight the improvement of our works.
Metrics. We evaluate the QA tasks and Misc by F1 Score [4] and Sequence Match Ratio [^5] https://docs.python.org/3/library/difflib.html for Retrieval, code tasks by exact match. Besides, we introduce fine-grained Progress Score $f$ from AgentBoard [54] for quantification of the multi-turn LLM agents’ performance which is denoted by:
$r_{t}=\max_{i,0\leq i\leq t}\frac{|\mathcal{G}\cap\mathcal{P}_{i}|}{|\mathcal{G}|}$ (9)
where an overall goal can be done by conjunction $g_{1}\land g_{2},\dots,\land g_{m}$ of atomic subgoal $\mathcal{G}=\{g_{1},g_{2},\dots,g_{m}\}$. While $\mathcal{P}_{i}=\{p_{1},p_{2},\dots,p_{m}\}$ express processed subgoals at $i^{th}$ step.
[…]
HotpotQA NarrativeQA Loogle-CR-mixup Loogle-MR-mixup Code.Run Code.Debug GraphWalks OpenAI-MRCR
Avg len 9151 18409 99.7k 95.4k 75.2k 114.7k 237.9k 223.5k
Max len 16k 65k 256k 256k 80k 200k 1m 1m
Task Type Multi-doc QA Single-doc QA Multi-hop QA Multi-hop QA Code Code Misc Retrieval
Eval Metric F1 F1 keyword-recall-based F1 keyword-recall-based F1 Accuracy Accuracy F1 Seq Match
4.2 Overall Results of XpandA
Table 2: Overall Performance evaluation of XpandA on benchmarks for Question-Answering, Code, Miscellaneous, and Retrieval, showing XpandA achieved optimal or leading Performance over full-context, RAG and previous agent-based works across multiple domains (Wilcoxon Signed-Rank Test: p<0.05, details in Appendix LABEL:apd:st).
Question Answering Code Tasks Misc. Retrieval
LLMs Method HotpotQA Narr.QA CR MIR Code.Debug Code.Run GraphWalks MRCR
Qwen2.5-7B-Inst Full-context 48.7 11.9 5.94 4.01 21.1 9.1 14.9 5.95
RAG 53.8 14.2 7.54 5.29 22.1 3.2 25.6 8.8
XpandA 67.3 20.7 15.1 12.6 24.8 21.2 29.8 16.3
LLama3.1-7B-Inst Full-context 53.2 12.3 8.3 7.41 20.2 7.7 14.9 2.09
RAG 55.1 15.7 11.8 10.2 25.4 1.4 15.5 12.9
XpandA 69.4 24.5 10.1 11.9 24.3 20.6 13.3 14.5
Qwen2.5-7B-Inst-1M Full-context 56.5 20.1 8.78 7.87 27.8 21.4 29.3 5.19
RAG 57.1 20.3 9.42 7.49 25.3 2.3 27.6 5.57
CoA 62.1 25.3 13.4 9.7 33.3 18.4 30.3 11.9
LongAgent 57.1 21.9 14.2 8.8 34.2 17.7 31.1 9.85
XpandA 64.5 26.4 17.4 13.9 40.5 25.8 32.5 15.3
Figure 2: F1 Score of Full-context, RAG and XpandA on all three backbone models in datasets LooGLE-MIR-mixup and LooGLE-CR-mixup of length varing from 16k to 256k. Xpanda achieves superiority on 5/6 datasets. Besides, Xpanda mitigates the performance decay in the case of input sequences of different lengths and reduces the gap between different generations of backbone models.
Question Answering. Table 2 shows the results of QA tasks on three backbone models. XpandA outperformed the full-context models on all seven datasets, with improvements of up to 18.6% on HotpotQA, 12.2% on NarrativeQA, 9.16% on LooGLE-CR-mixup, and 8.39% on LooGLE-MIR-mixup, respectively. Furthermore, XpandA surpassed RAG systems that utilize the same backbone model as their generator in performance. Notably, XpandA’s performance on QA tasks also exceeded that of previous agent-based methods, CoA and LongAgent, with gains of up to 5.1%.
Code Tasks, Retrieval & Misc. Table 2 displays the test results for these three classification tasks. Similarly, XpandA achieved optimal scores on the evaluation metrics for the majority of these tasks. However, it is noteworthy that RAG + Llama3.1-8B-Instruct delivered the best performance on the GraphWalks dataset, with XpandA ranking second. This outcome may be attributed to the fact that the text in GraphWalks is presented in a structured "node->node" format (where each ’node’ is a hexadecimal hash string). Such a structure potentially allows RAG to retrieve the intended textual information with greater precision compared to attention-based language models.
4.3 Comparison with RAG and Long Context LLMs
Table 2 shows that XpandA’s optimal performance in eight datasets is greater than the optimal performance of RAG on the corresponding datasets, with increases of $3.4\%~{}22.6\%$ respectively. It is noted in Figure 2 that Llama + RAG performs slightly better than Llama + XpandA, possibly due to Llama’s weaker ability to follow the multi-agent workflow instructions in XpandA, leading to a performance decrease. Additionally, RAG performs poorly in multi-hop tasks, likely because it cannot iteratively retrieve and reason about information step-by-step, a core mechanism enabling XpandA to handle such problems effectively. For tasks involving 16k to 1M tokens, we evaluate Qwen2.5-7B-Instruct-1M, a long-context language model (LCLM) supporting up to 1 million tokens, against models with smaller context windows (SCLMs, typically 128k tokens). Figure 2 shows that while LCLMs outperform SCLMs on 64k–256k token inputs, both exhibit performance degradation as sequences lengthen. XpandA mitigates this decay, maintaining stable F1 scores for 16k–256k inputs and enabling SCLMs to match LCLM performance on 256k-token tasks.
4.4 Comparison with Other Multi-Agent Frameworks
Table 3: Comparison with other agent-based works on GraphWalks and MRCR. XpandA significantly outperforms LongAgent and CoA in both success and progress rates on GraphWalks and MRCR, demonstrating superior efficiency in information acquisition and structured reasoning for multi-hop long-context problems.
LongAgent CoA XpandA
success progress/steps success progress/steps success progress/steps
Graphwalks Depth=2 53.2 63.2/6.4 55.6 71.8/3 62.4↑6.8 80.2↑8.4/3.9
Depth=4 21.4 27.7/8.5 6.25 21.4/3 38.2↑16.8 53.6↑25.9/6.5
Depth=8 6.7 10.2/8.9 2.3 11.4/3 19.4↑12.7 42.9↑31.5/7.7
MRCR 2 Needles 61.3 63.7/6.1 66.4 68.6/3 71.8↑5.4 77.8↑9.2/4.3
4 Needles 25.7 25.4/8.2 29.4 54.3/3 39.1↑9.7 68.3↑14.0/5.1
8 Needles 5.1 12.3/8.8 19.4 40.5/3 20.6↑1.2 44.8↑4.3/8.2
As shown in Table 3, XpandA outperforms LongAgent and CoA in success rate by up to 16.8% and progress rate by up to 32.7% on GraphWalks, and by 13.4% and 42.9% respectively on MRCR. It is noteworthy that, when compared to CoA, which lacks a replay mechanism, XpandA’s primary additional computational cost is the necessary overhead in operational steps due to its replay process. The higher progress rate reflects that XpandA more purposefully acquires information when traversing text chunks and more efficiently identifies the requisite information for each step in multi-hop problem-solving. The higher success rate signifies that XpandA develops a structured understanding of the collected information, thereby providing better answers rather than relying on superficial or decontextualized interpretations.
5 Analysis
5.1 Diverse Chunking Strategies Meet Context of Vastly Varying Length
Figure 3: Impact of different chunking strategies on Xpanda in dataset MRCR (2 Needle). Optimal static performance occurs when the unit chunk size is $[\frac{1}{16},\frac{1}{4}]$ of the input length, while dynamic chunking controls the proportion within this range and achieves 18% average improvement.
We analyze the impact of chunking strategies on XpandA’s performance, shown in Figure 3, via ablation experiments on the MRCR (2 Needles) benchmark (16k–1M sequences), evaluating retrieval success rate. Static chunking (2k–128k) shows that when the chunk size exceeds the input length, XpandA degenerates to a single-agent system, suffering full-context-like performance decay. Conversely, overly small chunks introduce excessive agents, collecting irrelevant noise and degrading accuracy. Optimal static performance occurs when unit chunk size is $[\frac{1}{16},\frac{1}{4}]$ of the input sequence length.
XpandA’s dynamic chunking adaptively divides the input sequence into four to eight equal parts which fall into the range, thus achieving integrated peak performance of static strategy. This improves the average retrieval success rate by 18% over the best static variant.
5.2 LLM Reasoning Scheme Ablation
(a) Reasoning scheme (F1 Score).
(b) Replay strategy (Time).
(c) Replay strategy (F1 Score).
Figure 4: Ablation studies on NarrativeQA. (a) Comparison of XpandA with alternative planning schemes. (b) Impact of different replay strategies on average inference time. (c) Impact of different replay strategies on F1 score. Selective replay in XpandA achieves a strong balance between performance and efficiency.
To further explore the effectiveness of the question-driven scheme in xpanda, we systematically remove and modify the component from XpandA to observe the performance changes. Plan-and-Solve [78] and Fact-and-Reflection [103] can help improve the LLM reasoning performance through the management of atomic plans or atomic facts, behaving just like the subquestion state tracking in XpandA, and thus can replace the planning scheme in XpandA without reconstruction of pipeline. We implement Directly Answer by setting the question tracker $P$,$T$ to null value and gathering all Explorers’ direct answers to be chosen by Decider. As shown in Figure 4(a), XpandA outperforms other reasoning schemes by 4–12% in F1 score of NarrativeQA, which shows that the inclusion of question-driven scheme in XpandA efficiently reinforces the multi-agent workflow by providing instruction that is more explicit to identify and explore the context and mitigates the inter-agent misalignment.
5.3 Selective Replay Matters in Question-driven Workflow
Figure 4(b), 4(c) illustrate the performance of XpandA on the GraphWalks dataset (context length 128k) in terms of runtime and F1 score, comparing its standard selective replay mechanism against three alternative modes: no-replay, random replay, and brute-force replay (the latter executed either unidirectionally or by alternating directions with each replay). In the random replay setup, the number of chunks processed by selective replay was first recorded, and then an equivalent number of unique chunks was randomly selected and replayed sequentially. The experimental results indicate that selective replay achieves a nearly identical F1 score to that of the best-performing brute-force bidirectional replay, while utilizing only 56.5% of the latter’s runtime. The superiority in F1 score is primarily attributed to selective replay’s fine-grained, atomic management of information states, which enables it to select chunks for replay that are most likely to uncover beneficial information and drive problem resolution. Furthermore, its enhanced time efficiency can also be ascribed to selective replay’s ability to determine if the currently collected information is sufficient to answer the question, thereby enabling early stopping.
5.4 Empirical Inference Time Analysis
Figure 5: XpandA scheme exhibits 1.5x speedup in inference time compared to SOTA long context language model on long texts.
We analyze the inference time of XpandA on 4×A100_80G GPUs (Figure 5), using Qwen2.5:7B-Instruct-1M as the backbone to avoid API-related latency. For comparison, we include Llama3.1-Instruct-1048k [62] in full-context mode. XpandA with dynamic chunking outperforms all baselines, achieving a 1.5× speedup over Qwen2.5-1M (with SOTA sparse attention optimization [91]) while maintaining linear complexity with fixed chunk sizes. In contrast, traditional full-context models (limited to 256k context on 4×A100) are significantly slower. Dynamic chunking also balances inference time and model-call latency better than static strategies (2k/8k) used in CoA [100] and LongAgent [101], which suffer from accumulated per-agent latency—especially critical for commercial API calls.
6 Conclusion
In this paper, we propose XpandA, a novel paradigm for multi-agent LLM workflow in long context language modeling through continuous questioning-answering. It is a plug-and-play, robust, interpretable, and efficient multi-agent system framework that introduces a dynamic chunking strategy and selective replay in its question-driven workflow to further enhance its performance. Extensive experiments show that XpandA achieves superior performance in various tasks with a wide span of context lengths (10k to 1M) compared to Full-context, RAG, and previous agent-based methods. It is worth noting that through XpandA expansion, SOTA (state-of-the-art) long context models have seen further improvements in performance (20% metrics improvement and 1.5x speedup). Analysis indicates that XpandA, through the multi-module coupling of its question-driven workflow, effectively alleviates the failure problem of multi-agent systems, performing better in long-text processing scenarios.
7 Limitations
Although XpandA provides an excellent paradigm for addressing the weak constraints and operational inefficiencies of Long Context Multi-Agent Systems in prior research, future directions can address the following limitations to further enhance its capabilities and efficiency. First, the scalability of XpandA on models with weak instruction-following abilities could be improved through in-context learning, simplifying workflow instructions, or exploring more efficient workflows. Second, the current experiments limit input sequence lengths to under 1M; future work could further investigate XpandA’s behavior and performance on longer sequences.
Appendix A Proof of Completeness
Hypothetical Premise: the LLM($I$,$\theta$) acting an agent can complete the problem decomposition and reasoning of a specified context chunk within k tokens context window.
Problem Definition: Reasoning based on information distributed in different chunks has a certain dependency sequence. If chunks are not processed in this sequence, it may lead to the breakage of the reasoning process. This sequential dependency between reasoning across different chunks can be reduced to a topological sorting problem of predecessor-successor relationships in an AOV (Activity On Vertex) network. In the scenario of long text reasoning, the dependency order of chunks can change along with the variation of the entities of concern, which can be expressed as:
$M=\begin{pmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,y}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,y}\\ \vdots&\vdots&\ddots&\vdots\\ m_{x,1}&m_{x,2}&\cdots&m_{x,y}\end{pmatrix}$ (10)
where $m_{x,y}$ denotes the the $y^{th}$ chunk should be dealt with on the ${m_{x,y}}^{th}$ one in the topological sequence. For example, $\{4,2,1,3\}$ means reasoning should be done one-by-one on chunk 3, 2, 4, 1 successively.
Explorer scans through the chunks in one direction and Decider proposes replay when reaching the last column. The next starting point is set at the chunk, denoted as $o^{th}$ chunk, where the last unanswered problem $u$ is proposed. The starting points of different rows are the same.
$o=\max_{i\in[1,x]}\left(\arg\max_{j}u_{i,j}\right)$ (11)
The Explorer accepts $m_{x,y}$=1 and then 2, and so on until all chunks are received. Specifically, in ($y-1$) rounds of scan direction reverse, the Explorer would accept all items in M.
For the first starting point of reverse:
$\forall i\in[1,x],\forall j>o,m_{i,j}\neq m_{i,o}+1$ (12)
$\forall i\in[1,x],\exists j_{0}\in[1,o],m_{i,j_{0}}=m_{i,o}+1$ (13)
The starting point of the second round of reverse can be determined by:
$o=\min_{i\in[1,x]}\left(\arg\min_{j}u_{i,j}\right)$ (14)
For the second starting point of reverse:
$\forall i\in[1,x],\forall j<o,m_{i,j}\neq m_{i,o}+1$ (15)
$\forall i\in[1,x],\exists j_{0}\in[o,y],m_{i,j_{0}}=m_{i,o}+1$ (16)
The two processes are mirrored,
$M^{mir}=\begin{pmatrix}m_{1,y}&m_{1,y-1}&\cdots&m_{1,1}\\ m_{2,y}&m_{2,y-1}&\cdots&m_{2,1}\\ \vdots&\vdots&\ddots&\vdots\\ m_{x,y}&m_{x,y-1}&\cdots&m_{x,1}\end{pmatrix}$ (17)
Augment the matrix by concatenate $M$ and $M^{mir}$,
$\widetilde{M}=\Big{[}M\Big{|}M^{\text{mir}}\Big{|}M\Big{|}M^{\text{mir}}\Big{|}\cdots\Big{]}_{x\times y^{2}}.$ (18)
Scan from left to right. For every y chunks, at least one chunk is accepted. So all items in M can be accepted in $y$ scan, namely $y-1$ replay. It’s not non-trivial, for:
$\forall x,DrA\ accept\ M.$ (19)
Appendix B Supplementary Experiments
(a) Full model
(b) RAG model
(c) XpandA
Figure 6: Dual-Needle-in-a-Haystack test. (a) Full model. (b) RAG model. (c) XpandA.
The Dual-Needle-in-a-Haystack test is designed to rigorously evaluate a model’s capability to concurrently identify, associate, and utilize two distinct pieces of information ("needles") embedded within a voluminous and distractor-rich textual corpus ("haystack"). This task extends beyond single-needle retrieval by imposing a more significant challenge on the model’s long-range dependency understanding, information integration, and sustained attention across intensive contexts. This appendix presents a detailed performance analysis of the Full-Context, RAG, and XpandA methodologies on this task.
The "Length of Haystack", represented on the X-axis of Figure 6, was varied across discrete token counts from 1k to 128k. Additionally, two distinct "needles" were synthetically generated for each trial, whose depth distribution is represented on the Y-axis of the heatmaps. For a given trial, the two needles were semantically independent but both were required to answer the query. The "Depth Percent of Needles", shown on the Y-axis of the heatmaps, indicates the predefined contextual zones where the needles were inserted.
Figure 6(a) presents the performance heatmap for the Full-Context method. The overall average accuracy achieved was 52.58%. The general trend of degrading accuracy is observed as the haystack length increases beyond approximately 64k tokens, indicating the increasing difficulty in processing, retaining, and associating information from both needles when the context window is saturated. Successfully retrieving and associating two needles becomes more challenging when one or both are deeply embedded and when their relative distance is large within an extensive context.
The RAG method in Figure 6(b) achieved an overall average accuracy of 64.01%. As the heatmap indicates, performance is sensitive to haystack length. While RAG’s retrieval mechanism improves overall accuracy compared to the Full-Context method (Figure 6(a)), it exhibits limitations in effectively locating and leveraging information in increasingly longer documents. Its effectiveness negatively correlates with document length, suggesting potential retrieval challenges in complex long-context scenarios.
Our proposed multi-agent framework, XpandA in Figure 6(c), achieved an overall average accuracy of 72.53%, representing a significant improvement over the RAG (64.01%) and Full-Context (52.58%) methods on this task. The heatmap provides compelling visual evidence of XpandA’s enhanced robustness, effectively mitigating the substantial performance degradation typically observed in baseline approaches, particularly under challenging conditions such as extended context lengths, deep needle locations, and considerable relative distances between needles.